### v2

# for Dynamic Neural Radiance Fields

Abstract
The accurate reconstruction of dynamic scenes with neural radiance fields is significantly dependent on the estimation of camera poses. Widely used structure-from-motion pipelines encounter difficulties in accurately tracking the camera trajectory when faced with separate dynamics of the scene content and the camera movement. To address this challenge, we propose DynaMoN. DynaMoN utilizes semantic segmentation and generic motion masks to handle dynamic content for initial camera pose estimation and statics-focused ray sampling for fast and accurate novel-view synthesis. Our novel iterative learning scheme switches between training the NeRF and updating the pose parameters for an improved reconstruction and trajectory estimation quality. The proposed pipeline shows significant acceleration of the training process. We extensively evaluate our approach on two real-world dynamic datasets, the TUM RGB-D and the BONN RGB-D Dynamic dataset. DynaMoN improves over the state-of-the-art both in terms of reconstruction quality and trajectory accuracy. We plan to make our code public to enhance research in this area.
I Introduction
Enabling novel view synthesis on dynamic scenes often requires multi-camera setups [1]. However, everyday dynamic scenes are often captured by one single moving camera, restricting the field of view [1]. In Simultaneous Localization and Mapping (SLAM) pipelines, multiple scene representations can be used. Commonly used explicit voxel and surfel representations are complemented by implicit truncated signed distance fields (TSDF) or neural radiance fields (NeRF). The latter allows the rendering of photorealistic novel views from new, reasonable camera positions [2, 3]. While many are restricted to static scenes, dynamic NeRF [4, 5, 6, 7] allows novel view synthesis for dynamic scenes.
To enable novel view synthesis using NeRF, accurate camera poses, usually retrieved via COLMAP [8] or the Apple ARKit, are essential. The structure-from-motion (SfM)-based approach often demands hours of computation to estimate reasonable camera poses and is challenged by large-scale dynamic scenes. To overcome the limitation of slow SfM approaches, recent works combined SLAM and NeRF [9, 10, 11, 2]. SLAM provides faster results for the camera trajectory than classic SfM. However, by nature, scenes are more often dynamic than static. Existing NeRF-SLAM [9, 10, 11, 2] methods assume a static scene which poses difficulties when dynamic scene content is present as shown in Fig 1.
Figure 1: NeRF-SLAM approaches rely on static scenes (left). Approaches such as InstantNGP [3] can be used to mask out dynamic scene content [2, 11] (left). Still, when using the masking, the scene representation lacks in quality (left). Fully dynamic NeRF provides an implicit 4D (3D+time) representation; however, it relies on offline SfM which can suffer in the presence of considerable motion (center). We consider dynamics via our motion-based segmentation (MS) and semantic segmentation (SS) already in the camera localization (right) and further during novel view synthesis. This enables a more robust camera tracking and novel view synthesis with a higher quality (ours, right).
In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach which can handle highly dynamic scenes in camera pose retrieval and dynamic scene representation. We utilize semantic segmentation and generic motion masks during our camera localization process to enable robust tracking of the camera path in a dynamic environment. These masks are further utilized for ray sampling to train an efficient dynamic NeRF, based on HexPlane [4], that refines the initial camera poses, leading to improved reconstruction quality in scenes with strong motions.
We conduct experiments on two challenging datasets, namely the dynamic subset of the TUM RGB-D [12] and the BONN RGB-D Dynamic [13] dataset. The considered masking in DynaMoN allows us to outperform existing approaches in terms of the absolute trajectory error and novel view synthesis results on these datasets.
II Related Works
As DynaMoN is a combination of dynamic NeRF with fast and robust motion-aware camera localization, we group related work in topics for camera localization and their employment for novel view synthesis and neural representation for dynamic scenes.
II-A Camera Localization and Scene Dynamics
SfM and SLAM [14, 15, 16, 17] are arguably the two most common approaches for robust camera localization. More specifically, SLAM often targets real-time localization of a camera coupled with mapping. In addition to visual information, some approaches can utilize complementary sensory signals such as depth images and IMU data to further improve the accuracy. A common branch of work employs ORB descriptors to build a 3D map and track the camera [14, 17, 15].
In addition to classical or partially learnable SLAM approaches, Teed and Deng [16] introduce an end-to-end learnable method. Their approach, termed DROID-SLAM, consists of a dense bundle adjustment layer which enables recurrent iterative updates of camera poses and pixel-wise depth.
Typical camera localization approaches rely on a static scene assumption. However, natural scenes are often dynamic [18, 19, 20, 21]. To address this, Yu et al. [18] introduce a dynamic SLAM approach building upon ORB-SLAM2 [17]. Their DS-SLAM integrates a semantic segmentation network and a moving consistency check in the tracking process. Rünz et al. [22] enrich RGB-D SLAM with object-awareness and semantic segmentation capabilities while DynaSLAM [23] combines geometric and deep learning-based segmentation to mask dynamic objects in an attempt to ignore them for mapping. Others split the scene into dynamic and static parts. Dai et al. [20] build upon ORB-SLAM2 [17] and introduce a semantic segmentation using point correlation for this split and Ye et al. [21] divide the scene representation in a static flow field and a dynamic flow field.
Following this idea, we also split the scene into static parts and dynamic content using predicted segmentation masks.
II-B Camera Localization and NeRF
With the retrieved camera poses, a captured environment can be represented in 3D. Aside from explicit maps, an implicit scene representation can be established using a multilayer perceptron (MLP) [24]. Mildenhall et al. [24] introduce the use of an MLP as a neural representation for novel view synthesis, denoted as NeRF. The implicit representation is regressed by a photometric comparison of rendered views with input images. This allows the synthesis of photorealistic images from novel viewpoints at the cost of the requirement for accurate camera poses during training, typically captured by computationally intensive SfM with COLMAP [8]. An efficient alternative to SfM is the use of SLAM to retrieve camera poses.
In this regard, Sucar et al. [25] propose the usage of an MLP as scene representation in a SLAM framework and Zhu et al. [9] arrange multiple NeRFs in a hierarchical feature grid to learn both camera pose and scene representation. NICER-SLAM [10] introduces a similar approach for RGB input with estimated depth. Others purely rely on RGB and combine SLAM and NeRF [2, 11] utilizing hash-based InstantNGP [3] for scene representation. NeRF-SLAM [2] adapts the initial Droid-SLAM [16] framework and Orbeez-SLAM [11] leverages visual odometry [17].
To incorporate efficient pose estimation, we also rely on design choices of recent SLAM frameworks; however, we combine them with computationally efficient proposals [4] for neural radiance fields.
II-C Neural Representation of Dynamic Scenes
Our world is inherently dynamic which violates commonly used static scene assumptions [9, 10, 25, 11, 2, 3, 24]. To represent dynamics in a NeRF [26, 27], the use of grid structures became popular [4, 5]. HexPlane [4] utilizes a 4D space-time grid divided into six feature planes. The feature vector is a 4D point in space-time projected onto each feature plane. Similarly, Tensor4d [5] builds upon a 4D tensor field using time-aware volumes projected onto nine 2D planes. TiNeuVox [7] represents dynamic scenes using time-aware voxel features to enable faster training.
Optimizing for both camera poses and view synthesis can improve results in static scenes [28, 29, 30, 31, 32]. Liu et al. [33] therefore introduce such an optimization for a dynamic environment. To model this, they use one static and one dynamic NeRF.
We, in contrast, split the problem into differentiation of static and dynamic content already for camera pose estimation. The dynamic environment is then represented by a dynamic, 4D NeRF calculated upon the camera poses from the static scene components, which is further utilized to refine these poses.
III Method
We consider our input to be a video captured in a dynamic environment using a single monocular moving camera without pose knowledge. This leads to two objectives: first, estimating the camera pose, and second, reconstructing the scene which may comprise some deformation. We leverage masks to separate scene dynamics from static background for the former and utilize a 4D spatio-temporal neural representation for the latter. This allows to optimize for novel view synthesis while refining initial camera poses.
Figure 2: DynaMoN architecture. DynaMoN retrieves consecutive frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$). Combining these masks (OR), we enable a motion-aware, fast and robust camera pose estimation. Based on the predicted camera poses $G$, we produce a time-dependent output ($\Theta$) using dynamic NeRF. During NeRF training, we utilize the previously computed masks to refine the camera poses (dashed lines). Dashed lines represent iterative updates on NeRF and pose parameters. We highlight our approach in green.
III-A Robust Camera Localization in Dynamic Environments
Novel view synthesis methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation. This assumes a static scene. To cope with dynamics, it is possible to add binary masks to remove non-static regions. However, the high computation time remains a drawback. In our approach, we target both robustness against dynamics and a more efficient solution for pose estimation. Our underlying architecture is based on DROID-SLAM [16]. While it incorporates elements to cope with a certain level of scene dynamics, we found it can be improved by eliminating the contribution of the non-static regions for camera tracking. We do this by generating binary dynamics masks for each frame, which are combinations of semantic ($M_{SS}$) and motions masks ($M_{MS}$).
The semantic mask follows the paradigm that in real videos, the most prominent source of motion is often an object of known class. We therefore define $M_{SS}=\mathbb{1}(x>0)\circ\sum_{i}M_{SS}(i)$ for all defined semantic classes $i\in S$ of the set $S$ with the indicator function being $1$ for positive values. Additional moving objects are covered through a motion mask $M_{MS}$. One example is depicted by the moving box in Fig 2. Following DytanVO [34], our motion masks are generated using a CNN with input of previous image, forward optical flow and estimated camera motion to iteratively refine the mask along with pose updates. Instead of random initialization, we begin with the motion mask of the previous frame to improve stability and lower the number of iterations in our refinement. We combine both masks as $M=\mathbb{1}(x>0)\left(M_{SS}+M_{MS}\right)$.
III-B Neural Dynamic Scenes and Pose Refinement
Our scene representation builds on one of the recent dynamic NeRF models, HexPlane [4]. Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [35] and constructs a 4D representation with three spatial and one temporal axis to create feature planes with their pairwise combinations to be the input of their tiny MLP. The grid size of each axis is governed in a coarse-to-fine fashion along the training to improve convergence and naturally regularizes the local neighborhoods. For optimization, it follows the original NeRF’s ray casting formulation and applies a pixel-wise photometric loss between the rendered RGB values $C$ and the ground truth $\hat{C}$, supported by a total variation regularization as:
$\mathcal{L}=\frac{1}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}\|C(\mathbf{r})-\hat{C}(\mathbf{r})\|_{2}^{2}+\lambda_{\text{TV}}\mathcal{L}_{\text{TV}}.$ (1)
While $\mathcal{R}$ represents the sampled rays, $\mathcal{L}_{\text{TV}}$ and $\lambda_{\text{TV}}$ depict the total variation regularization and its weight.
TABLE I: RMSE of the translational ATE [m] on TUM RGB-D [12]. The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes. The - denotes that no results were reported for these sequences.
Sequences RGB-D Monocular
DVO-SLAM  [36] ORB-SLAM2 [17] PointCorr [20] DynaPix-D [37] DytanVO [34] DeFlowSLAM [21] DROID-SLAM [16] OURS (MS&SS) OURS (MS)
fr2/desk-person 0.104 0.006 0.008 - 1.166 0.013 0.006 0.007 0.006
fr3/sitting-static 0.012 0.008 0.010 - 0.016 0.007 0.005 0.005 0.005
fr3/sitting-xyz 0.242 0.010 0.009 - 0.260 0.015 0.009 0.010 0.009
fr3/sitting-rpy 0.176 0.025 0.023 - 0.046 0.027 0.022 0.024 0.021
fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.310 0.025 0.014 0.023 0.019
fr3/walking-static 0.752 0.408 0.011 0.007 0.021 0.007 0.012 0.007 0.014
fr3/walking-xyz 1.383 0.722 0.087 0.014 0.028 0.018 0.016 0.014 0.014
fr3/walking-rpy 1.292 0.805 0.161 0.123 0.155 0.057 0.040 0.031 0.039
fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.385 0.420 0.022 0.019 0.020
Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.016/0.039
III-C Pose Refinement and Training Strategy
We utilize eqn. (1) as the objective function for training both the NeRF and refining the camera poses. For the latter, we employ an alternating training strategy in which every 5,000 iterations the parameters to be optimized are switched between the radiance field and the pose parameters. To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, to be elements of the static regions designated by the dynamics masks. Furthermore, we also use the same masks to ensure that the total variation regularization is imposed only on the static components. As the amount of non-dynamic pixels can vary from image to image, we accordingly scale the learning rate by the amount of static pixels in the image.
Figure 3: RMSE of the rotational ATE [°] on the TUM RGB-D dataset (left) and the BONN RGB-D Dynamic dataset (right). The median is shown in orange and the average in dotted lines.
IV Evaluation
DynaMoN is assessed on two challenging datasets, examining camera localization and novel view synthesis quality. To evaluate our approach, we consider two aspects. First, we consider the translational and rotational absolute trajectory error (ATE) [38] using root mean square error (RMSE) in meters and degrees. Second, we analyze the rendering quality of the novel dynamic views by reporting PSNR and SSIM [39] values.
TABLE II: RMSE of the translational ATE [m] on BONN RGB-D [13]. RGB-D results are reported by Palazzolo et al. [13]. For DROID-SLAM and ours, we calculated the results.
Sequences RGB-D Monocular
ReFusion [13] StaticFusion [40] DynaSLAM (G) [23] DynaSLAM (N+G) [23] DROID-SLAM [16] OURS (MS&SS) OURS (MS)
balloon 0.175 0.233 0.050 0.030 0.075 0.028 0.068
balloon2 0.254 0.293 0.142 0.029 0.041 0.027 0.038
balloon_tracking 0.302 0.221 0.156 0.049 0.035 0.034 0.036
balloon_tracking2 0.322 0.366 0.192 0.035 0.026 0.032 0.028
crowd 0.204 3.586 1.065 0.016 0.052 0.035 0.061
crowd2 0.155 0.215 1.217 0.031 0.065 0.028 0.056
crowd3 0.137 0.168 0.835 0.038 0.046 0.032 0.055
kidnapping_box 0.148 0.336 0.026 0.029 0.020 0.021 0.020
kidnapping_box2 0.161 0.263 0.033 0.035 0.017 0.017 0.017
moving_no_box 0.071 0.141 0.317 0.232 0.023 0.013 0.014
moving_no_box2 0.179 0.364 0.052 0.039 0.040 0.027 0.026
moving_o_box 0.343 0.331 0.544 0.044 0.177 0.152 0.167
moving_o_box2 0.528 0.309 0.589 0.263 0.236 0.175 0.176
person_tracking 0.289 0.484 0.714 0.061 0.043 0.148 0.024
person_tracking2 0.463 0.626 0.817 0.078 0.054 0.022 0.035
placing_no_box 0.106 0.125 0.645 0.575 0.078 0.021 0.018
placing_no_box2 0.141 0.177 0.027 0.021 0.030 0.020 0.020
placing_no_box3 0.174 0.256 0.327 0.058 0.025 0.022 0.022
placing_o_box 0.571 0.330 0.267 0.255 0.127 0.172 0.117
removing_no_box 0.041 0.136 0.016 0.016 0.016 0.015 0.015
removing_no_box2 0.111 0.129 0.022 0.021 0.020 0.019 0.021
removing_o_box 0.222 0.334 0.362 0.291 0.189 0.177 0.181
synchronous 0.441 0.446 0.977 0.015 0.006 0.007 0.006
synchronous2 0.022 0.027 0.887 0.009 0.012 0.006 0.011
Mean/Max 0.232/0.571 0.412/3.586 0.428/1.217 0.095/0.575 0.061/0.236 0.052/0.177 0.051/0.181
To assess DynaMoN, we compare both camera localization and novel view synthesis part with state-of-the-art approaches on two challenging datasets. For the novel view synthesis results, we follow the usual splitting using every 8th frame of the TUM RGB-D [12] dataset and BONN RGB-D Dynamic [13] dataset for testing.
IV-A Implementation Details
We implement our approach in Python using PyTorch. For semantic masking, we use DeepLabV3 [41] with a ResNet50-backbone [42] trained on the large-scale COCO dataset [43] and use human masks. For motion masking, we use the trained model of DytanVO [34]. In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence. For cases where the camera motion is higher than in the training sequences of DytanVO, the motion segmentation module results in an increase of false positives; as a counter-measure, we introduce a threshold for the number of pixels that contain motion. Consequently, masks exceeding this threshold are discarded. This ensures that there are always enough pixels for the dense bundle adjustment.
While we use the reported image sizes for the novel view synthesis experiments, we use half of the resolution for the camera localization to be consistent with existent trajectory evaluations of DROID-SLAM [16].
Regarding the loss function, we set $\lambda_{TV}$ initially to 0.005 and the number of sampled rays $\lvert\mathcal{R}\rvert$ per batch to 1024.
To compare our camera pose retrieval with the classic way of using COLMAP, we follow the implementation of InstantNGP [3] to retrieve the COLMAP ground truth. Besides the COLMAP comparison, we compare our approach with RoDynRF [33]. For the TUM RGB-D dataset, we train their approach on an NVIDIA RTX 3090 with 24GB of VRAM. We follow their training guideline and do not employ prior pose initialization.
TABLE III: Novel view synthesis results on TUM RGB-D [12]. The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes. The - denotes that the camera pose could not be regressed.
[…]
RoDynRF
HexPlane
HexPlane
+ DynaMoN
Backbone (Ours)
DynaMoN
fr2/desk-person 6.05 / 0.127 - 23.04 / 0.680 24.78 / 0.724
fr3/sitting-static 9.67 / 0.123 - 28.55 / 0.899 28.95 / 0.903
fr3/sitting-xyz 10.10 / 0.150 19.98 / 0.589 25.90 / 0.848 26.15 / 0.850
fr3/sitting-rpy 8.35 / 0.121 16.63 / 0.472 25.76 / 0.843 26.40 / 0.853
fr3/sitting-halfsphere 7.46 / 0.157 24.25 / 0.793 24.05 / 0.792 24.60 / 0.799
fr3/walking-static 10.01 / 0.159 20.27 / 0.650 25.84 / 0.849 26.25 / 0.853
fr3/walking-xyz 10.02 / 0.171 - 24.05 / 0.797 24.34 / 0.796
fr3/walking-rpy 8.81 / 0.201 - 24.18 / 0.790 24.70 / 0.796
fr3/walking-halfsphere 8.79 / 0.205 22.01 / 0.692 24.05 / 0.781 24.27 / 0.779
Mean 8.84 / 0.157 - 25.05 / 0.809 25.60 / 0.819
TABLE IV: Novel view synthesis results on BONN RGB-D Dynamic [13]. The - denotes that the camera pose could not be regressed.
[…]
HexPlane
HexPlane +
DynaMoN
Backbone (Ours)
DynaMoN
balloon - 29.83 / 0.868 29.84 / 0.866
balloon2 - 27.61 / 0.841 28.31 / 0.847
balloon_tracking - 29.27 / 0.847 29.55 / 0.846
balloon_tracking2 - 29.68 / 0.860 30.56 / 0.862
crowd - 27.92 / 0.831 28.40 / 0.829
crowd2 - 26.68 / 0.825 27.97 / 0.827
crowd3 - 26.57 / 0.815 27.40 / 0.827
kidnapping_box - 30.43 / 0.870 30.59 / 0.869
kidnapping_box2 30.87 / 0.881 30.28 / 0.860 30.18 / 0.857
moving_no_box 30.67 / 0.881 30.32 / 0.866 30.36 / 0.862
moving_no_box2 31.14 / 0.890 30.57 / 0.874 30.62 / 0.870
moving_o_box 29.21 / 0.889 30.26 / 0.884 31.08 / 0.888
moving_o_box2 - 29.82 / 0.875 29.40 / 0.875
person_tracking 26.57 / 0.808 27.78 / 0.843 28.71 / 0.845
person_tracking2 - 27.47 / 0.837 27.80 / 0.833
placing_no_box 25.73 / 0.780 30.94 / 0.869 31.23 / 0.870
placing_no_box2 - 31.18 / 0.885 31.45 / 0.884
placing_no_box3 - 30.79 / 0.879 30.83 / 0.875
placing_o_box - 30.42 / 0.886 31.12 / 0.890
removing_no_box - 30.66 / 0.879 30.85 / 0.877
removing_no_box2 31.13 / 0.887 30.73 / 0.871 30.86 / 0.870
removing_o_box 31.13 / 0.887 30.30 / 0.880 30.76 / 0.882
synchronous - 28.42 / 0.835 28.93 / 0.834
synchronous2 - 28.85 / 0.887 30.18 / 0.891
Mean - 29.45 / 0.861 29.87 / 0.861
Figure 4: Qualitative Results for Novel View Synthesis on TUM RGB-D dataset (top) and BONN RGB-D dynamic dataset (bottom). Shown are original Ground Truth images (left) and synthesized views from COLMAP-based dynamic NeRF (2nd column) with improved camera poses from our backbone (3rd column) only, and our full model (right).
IV-B Camera Localization Quality
Our results on all datasets show a lower translational trajectory error compared to state-of-the-art approaches, see Table I and Table II. We compare RGB-D-based and monocular SLAM approaches with our proposed camera retrieval method using only motion masks (MS) as well as using motion and semantic segmentation masks in combination (MS&SS). Looking at the mean value of the translational trajectory error, our approach performs similarly to DROID-SLAM on TUM RGB-D. However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world sequences. Moreover, comparing the rotational and translational ATE, see Figure LABEL:fig:atel, our MS&SS achieves lower values than DROID-SLAM. On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error.
TABLE V: Analysis of the influence of different pose retrieval steps to PSNR and SSIM in the Bonn RGB-D Dynamics crowd2 scene with sparse input trajectory.
Approach DROID-SLAM + HexPlane
Our initial poses + HexPlane
DynaMoN
PSNR $\uparrow$
[…]
16.46 / 0.661 17.07 / 0.660 17.31 / 0.659
IV-C Novel View Synthesis Quality
To evaluate our approach from a NeRF perspective we consider two aspects. First, the traditional way of generating camera ground truth when using NeRF, which is the use of COLMAP. Thus, we compare DynaMoN with HexPlane [4] using camera poses from COLMAP. Second, we highlight the influence of the iterative pose refinement by training both DynaMoN and HexPlane with our initially localized poses. Third, we compare our approach with RoDynRF [33] which regresses the camera pose along with a dynamic NeRF.
The results for the TUM RGB-D dataset are denoted in Table III. DynaMoN leads to a higher novel view synthesis quality compared to HexPlane without refined poses, RoDynRF, and when using COLMAP for the camera poses. Qualitative examples of our renderings are shown in Fig. 4.
On the BONN RGB-D dataset, COLMAP is challenged on several sequences (cf. Tab. IV). DynaMoN can retrieve camera poses for all scenes (Tab. II) and performs well on novel view synthesis for all sequences (Tab. IV). In addition, the PSNR is increased by the alternating training scheme. However, for the cases where COLMAP succeeds in the pose estimation, the camera poses from SfM also lead to decent novel view synthesis results. Examples of renderings can be seen in Fig. 4.
IV-D Robustness on Sparse Trajectories
Finally, we evaluate the robustness of different parts of our pose retrieval when faced with a dynamic and sparse trajectory. Using only every 10th frame of the crowd2 sequence, we show how the DROID-SLAM poses compare to the ones from the dynamics-masked camera pose retrieval and our neural camera pose refinement in terms of novel view rendering quality. Looking at the PSNR values in Table V, it can be noted that each of our proposed components incrementally increases the reconstruction quality. Thus, when confronted with a sparse trajectory of a highly dynamic scene, DynaMoN outperforms pure DROID-SLAM making it more robust to realistic, real-world scenarios with imperfect camera recordings.
V Discussion
In comparison to traditional sparse reconstructions, our approach addresses camera localization and novel view synthesis in combination. While this leads to more pleasing 3D visualizations and novel view synthesis results, the computing time is higher compared to traditional camera localization and 3D representation methods.
For camera pose estimation, our approach shows an improved performance compared to the state-of-the-art. Only using our motion masks already provides an improved performance. Moreover, for scenes with large motions, the combination with semantic segmentation masks provides significantly improved camera localization results. The results presented in Table III indicate that RoDynRF [33], despite being a state-of-the-art method that integrates combined camera regression with novel view synthesis, faces significant challenges in handling highly dynamic camera movements and dynamic scenes. In contrast, our method demonstrates superior robustness and yields improved outcomes when compared to both RoDynRF [33] and HexPlane [4] employing COLMAP poses on the demanding TUM RGB-D dataset. Furthermore, our iterative pose refinement process enhances robustness, as evidenced in Table V, and achieves better performance in novel view synthesis on both datasets.
VI Conclusion
We present DynaMoN, a motion-aware fast and robust camera localization approach for novel view synthesis. DynaMoN can handle not only the motion of known objects using semantic segmentation masks but also that of unknown objects using a motion segmentation mask. Furthermore, it retrieves camera poses faster and is more robust to scene dynamics in comparison to traditional SfM approaches, enabling a more accurate 4D scene representation. In addition, DynaMoN is trained with a novel iterative training scheme that refines NeRF parameters and the initially provided poses in an alternating fashion. Compared to the state-of-the-art, DynaMoN outperforms other dynamic camera localization approaches and shows better results for novel view synthesis on scenes with high camera motion.
Acknowledgment
We thank d.hip for providing a campus stipend.